# HuggingGraph: Understanding the Supply Chain of LLM Ecosystem

### v2

@@ 4.6. RQ #6: Dynamic Update Evaluation
This research question aims to understand the dynamic update of the LLM supply chain. Following the method discussed in Section 3.4, we perform a daily-based data collection by capturing how many nodes and edges are added and deleted each day.
@@
Figure 5 illustrates the sum of daily added and deleted of six key node categories (base, fine-tuned, adapters, quantized, merge variants, and datasets) from June 25 to July 15, 2025. We make three interesting observations. (i) The daily dynamic update is significant. That is, an average of 4,622 models are changing every day, including $\sim$3,843 model additions and $\sim$779 deletions. In addition, about 1,538 datasets are changing each day, containing $\sim$1,305 dataset additions and $\sim$233 deletions. An addition occurs when a new model or dataset is uploaded to the Hugging Face platform. This includes base models, task-specific variants, and new training datasets. A deletion refers to the removal of such nodes, often due to licensing issues, privacy concerns, or contributor decisions, such as replacing outdated models or withdrawing low-quality or sensitive datasets.
@@
(ii) Fine-tuned models dominate daily activity, averaging over 2,988 changes per day, followed by consistent contributions from adapters ($\sim$1,224/day) and datasets ($\sim$1,539/day). Noticeable spikes, such as on July 7 and July 9, align with major events like the Mistral-Fusion-v3 fine-tuning wave and dataset updates such as HFTime2025-News. (iii) Furthermore, adapter uploads peaked at 1,249 on June 28, while quantized variants reached 381 on July 9, driven by releases such as QWin-GGUF-7B. These patterns demonstrate HuggingGraph’s ability to capture evolving supply chain dynamics at a fine-grained level.
@@
Finding #6: The LLM supply chain exhibits continuous and high-volume daily changes, driven by frequent additions and deletions of models and datasets. This reflects a rapidly evolving and highly dynamic ecosystem shaped by active contributor behavior.
4.7. RQ #7: Generability to Other AI Platforms
@@
To validate HuggingGraph’s generalizability beyond Hugging Face, we applied our pipeline to another AI platform, Kaggle (Bojer and Meldgaard, 2021). As of July 25, 2025, Kaggle hosts 470 base and task-specific models with 3,146 variants and approximately 502K datasets. Using Kaggle’s kernel and dataset APIs, we collected 2,640 models and 105,867 datasets. This significant gap is due to lack of the models’ and datasets’ metadata. Of the datasets retrieved ($\sim$105K), only 137 datasets were included in our graph, as most lacked standardized documentation or traceable links to models. Many are standalone, poorly described, or lack contextual information, a challenge also observed on Hugging Face.
@@
We follow the same way to construct a heterogeneous graph consisting of 2,777 nodes, which include 2,640 model nodes, comprising 59 base models, 2,410 fine-tuned, 171 quantization models, and 137 datasets. The graph contains a total of 3,990 edges, on which we observed seven types of edges, (i) base model $\rightarrow$ fine-tuned model (467 edges), (i) base model $\rightarrow$ quantization model (62 edges), (iii) fine-tuned model $\rightarrow$ fine-tuned model (1,696 edges), (iv) fine-tuned model $\rightarrow$ quantized model (107 edges), (v) quantized model $\rightarrow$ quantized model (1 edges), (vi) dataset $\rightarrow$ fine-tuned model (1,614 edges), and (vii) dataset $\rightarrow$ quantization model (43 edges).
@@
We observed that the average degree is 1.44, indicating that the graph is sparse. We made two interesting observations. (i) The degree distribution is heavy-tailed and skewed: out of 2,777 total nodes, 2,305 nodes ($\approx$83%) have a total degree of 1. Most nodes have low degrees, while a few highly connected hubs dominate the graph. For example, in-degrees range from 0 to 4 (with NaN peaking at 4), and tensorflow/mobilenet-v1 has the highest out-degree of 64, followed by google/nnlm with 56. (ii) Furthermore, the graph contains 448 weakly connected components (WCCs), reflecting high fragmentation. However, the largest WCC includes 65 nodes, suggesting the presence of a moderately sized core subgraph.
@@
Finding #7: The resulting graph exhibits structural properties consistent with our Hugging Face analysis, including a heavy-tailed degree distribution, sparse connectivity, and strong modular fragmentation, demonstrating the robustness and generalizability of our pipeline across platforms despite metadata limitations.
@@ 5. Use Case
HuggingGraph presents a technique to analyze the supply chain of the LLM ecosystem. The proposed graph can be used for various applications, e.g., auditing provenance, identifying biases, and revealing trends like quantized model scarcity, aiding researchers and platform maintainers. We discuss the following two use cases.
@@
Use case #1: Tracing lineage and dependencies in the LLM supply chain. In the LLM ecosystem, models are frequently built upon others through fine-tuning, adapter training, or quantization, forming complex chains of dependencies. However, when these relationships are not explicitly visible, it becomes difficult to verify where a model comes from, whether it inherits bias from upstream datasets, or if it complies with licensing constraints. HuggingGraph can be used to address this challenge by constructing the supply chain of models and datasets, uncovering both direct and derived dependencies, even when they are not formally documented. For example, it can trace how the model Meta-llama indirectly relies on a dataset like Wikimedia via Chatgpt-prompt (Figure 2). This transparency supports developers, auditors, and policymakers in validating provenance, detecting risks, and enabling trustworthy AI.
@@
Use case #2: Identifying critical nodes and structural vulnerabilities. In the LLM ecosystem, certain models (e.g., gemma-2b) and datasets (e.g., The Pile) are reused so frequently that they become critical structural hubs, where failure or removal of them could disrupt numerous downstream dependencies. These hidden single points of failure are difficult to detect without a comprehensive view of resource interconnections. HuggingGraph can be used to address this by modeling the supply chain as a graph and analyzing node connectivity to surface highly reused models and datasets with significant inbound or outbound links. This visibility enables maintainers to safeguard vital assets and helps developers mitigate the risk of overreliance on fragile or under-maintained components.
6. Related Work
This section aims to provide in-depth coverage of the current research landscape concerning the opportunities within the LLM supply chain, mainly emphasizing the different dimensions of the relationship between the models and datasets, focusing on the model and dataset hosting platforms (Sindhu et al., 2024; Hadi et al., 2024).
LLM supply chain perspectives in AI. LLMs mark the revolutionary era in AI, with the edge passing through an exponential growth phase. A recent paper presents LLMs key components, including model infrastructure, lifecycle, and downstream applications (Wang et al., 2024). Another work showed that reusing these models has become widespread, which encourages the sharing and adaptation of base models on a larger scale (Jiang et al., 2023). An open-source AI ecosystem, such as Hugging Face, hosts a broad range of LLMs and datasets and, therefore, plays an important role in democratizing AI technologies (Riva et al., 2024). The base models are key ingredients in such an ecosystem (Xu et al., 2024b). They capture vast amounts of information from various datasets, allowing for effective task-specific variants like fine-tuning (Han et al., 2021). This trend underscores the significant impact of technological advancements on AI democratization and innovation (Cupać et al., 2024).
@@ 6. Related Work
Relationship analysis between LLM models and datasets. A recent study investigates the practical adaptation of base models to specific areas and tasks. Multitask fine-tuning has demonstrated the potential to enhance performance on target tasks with scarce labels (Xu et al., 2024a). In plant phenotyping, adapting vision-based models by techniques like adapter tuning and decoder tuning has shown results comparable to those of leading task-specific models (Chen et al., 2023). The Quadapter technique for language models tackles quantization difficulties by incorporating learnable parameters that scale activations channel-wise, mitigating overfitting during quantization-aware training (Park et al., 2022). These findings underscore the efficacy of various adaptation approaches in improving base model performance across areas.
From the above, we can see that models heavily depend on each other for fine-tuning, adaptation, quantization or merger. Building on these insights, two other works have indicated that serious security vulnerabilities may encumber the LLM supply chain to a few developers (Kathikar et al., 2023; Osborne et al., 2024). These factors may affect the diversity of innovation. The authors relate several challenges arising during the software engineering, security, and privacy of different relations and components involved in model creation and deployment (Huang et al., 2023; Wang et al., 2024). Moreover, the vulnerabilities can be transferable from one model to another model during the fine-tuning process (Hajipour et al., 2024).
@@ 7. Conclusion
This project studies the relationships between models and datasets in the LLM ecosystem, which are the central parts of the LLM supply chain. First, we design a methodology to systematically collect the supply chain information of LLMs. With that, we construct a large directed heterogeneous graph, having 402,654 nodes and 453,469 edges. Then, we perform different types of analysis and make multiple interesting findings.

### v3

@@ 4.5. RQ #5: Dynamic Update Evaluation
This research question aims to understand the dynamic update of the LLM supply chain. We perform a daily-based data collection by capturing the addition and deletion of nodes and edges each day.
@@
Figure 5 denotes the number of changed (additions and deletions) datasets and models from June 25 to July 15, 2025. We make three interesting observations. (i) The daily dynamic update is significant. That is, an average of 4,622 models are changing every day, including $\sim$3,843 model additions and $\sim$779 deletions. In addition, about 1,538 datasets are changing each day, containing $\sim$1,305 dataset additions and $\sim$233 deletions. An addition occurs when a new model or dataset is uploaded to the Hugging Face platform. This includes base models, task-specific variants, and new training datasets. A deletion refers to the removal of such nodes, often due to licensing issues, privacy concerns, or contributor decisions, such as replacing outdated models or withdrawing low-quality or sensitive datasets.
@@
(ii) Fine-tuned models dominate daily activity, averaging over 2,988 changes per day, followed by consistent contributions from adapters ($\sim$1,224/day) and datasets ($\sim$1,539/day). Noticeable spikes, such as on July 7 and July 9, align with major events like the Mistral-Fusion-v3 fine-tuning wave and dataset updates such as HFTime2025-News. (iii) Furthermore, adapter uploads peaked at 1,249 on June 28, while quantized variants reached 381 on July 9, driven by releases such as Qwen-GGUF-7B. These patterns demonstrate HuggingGraph’s ability to capture evolving supply chain dynamics at a fine-grained level.
@@
Finding #5: The LLM supply chain exhibits continuous and high-volume daily changes, driven by frequent additions and deletions of models and datasets. This reflects a rapidly evolving and highly dynamic ecosystem shaped by active contributor behavior.
4.6. RQ #6: Generalization to Other AI Platforms
@@
To validate HuggingGraph’s generalizability beyond Hugging Face, we applied our pipeline to another AI platform, Kaggle (Bojer and Meldgaard, 2021). As of July 25, 2025, Kaggle hosts 470 base models, 3,146 task-specific models, and $\sim$502K datasets. Using Kaggle’s kernel and dataset APIs, we collected 2,640 models and 105,867 datasets. This significant gap is due to a lack of models’ and datasets’ metadata. Of the datasets retrieved ($\sim$105K), only 137 datasets were included in our graph, as most lacked standardized documentation or traceable links to models. Many are standalone, poorly described, or lack contextual information, a challenge also observed on Hugging Face.
@@
We follow the same way to construct a heterogeneous graph consisting of 2,777 nodes, which include 2,640 model nodes, comprising 59 base models, 2,410 fine-tuned, 171 quantization models, and 137 dataset nodes. The graph contains a total of 3,990 edges, on which we observed seven types of edges, (i) base model $\rightarrow$ fine-tuned model (467 edges), (ii) base model $\rightarrow$ quantization model (62 edges), (iii) fine-tuned model $\rightarrow$ fine-tuned model (1,696 edges), (iv) fine-tuned model $\rightarrow$ quantized model (107 edges), (v) quantized model $\rightarrow$ quantized model (1 edge), (vi) dataset $\rightarrow$ fine-tuned model (1,614 edges), and (vii) dataset $\rightarrow$ quantization model (43 edges).
@@
We observed that the average degree is 1.44, indicating that the graph is sparse. We made two interesting observations. (i) The degree distribution is heavy-tailed and skewed: out of 2,777 total nodes, 2,305 nodes (83%) have a total degree of 1. Most nodes have low degrees, while a few highly connected hubs dominate the graph. For example, in-degrees range from 0 to 4, and tensorflow/mobilenet-v1 has the highest out-degree of 64, followed by google/nnlm with 56. (ii) Furthermore, the graph contains 448 WCCs, reflecting high fragmentation. However, the largest WCC includes 65 nodes, suggesting the presence of a moderately sized core subgraph.
@@
Finding #6: The resulting graph exhibits structural properties consistent with our Hugging Face analysis, including a heavy-tailed degree distribution, sparse connectivity, and strong modular fragmentation, demonstrating the robustness and generalizability of our pipeline across platforms despite metadata limitations.
@@ 5. Use Case
HuggingGraph presents a technique to analyze the supply chain of the LLM ecosystem, which can be used for various applications, e.g., auditing provenance, identifying biases, and revealing trends like quantized model scarcity. We discuss the two use cases.
@@
Use case #1: Tracing lineage and dependency in LLM supply chain. Models are frequently built upon others through fine-tuning, adapter training, or quantization, forming complex chains of dependencies. However, when these relationships are not explicitly visible, it becomes difficult to verify where a model comes from, whether it inherits bias from upstream datasets, or if it complies with licensing constraints. HuggingGraph can be used to address this challenge by constructing the supply chain of models and datasets, uncovering both direct and derived dependencies, even when they are not formally documented. For example, it can trace how the model Meta-llama indirectly relies on a dataset like Wikimedia via Chatgpt-prompt (Figure 2). This transparency supports developers, auditors, and policymakers in validating provenance, detecting risks, and enabling trustworthy AI.
@@
Use case #2: Identifying critical nodes and structural vulnerabilities. In the LLM ecosystem, certain models (e.g., gemma-2b) and datasets (e.g., The Pile) are reused so frequently that they become critical structural hubs, where their failure or removal could disrupt numerous downstream dependencies. These hidden single points of failure are difficult to detect without a comprehensive view of resource interconnections. HuggingGraph can be used to address this by modeling the supply chain as a graph and analyzing node connectivity to surface highly reused models and datasets with significant inbound or outbound links. This visibility enables maintainers to safeguard vital assets and helps developers mitigate the risk of overreliance on fragile or under-maintained components.
6. Related Work and Discussion
LLM supply chain perspectives in AI. LLMs are advancing across model infrastructure, lifecycle, and applications (Wang et al., 2024). Model reuse is widespread, promoting large-scale sharing and adaptation of base models (Jiang et al., 2023). Open-source ecosystems such as Hugging Face host diverse LLMs and datasets, democratizing AI (Riva et al., 2024). Base models (Xu et al., 2024b), trained on broad datasets, enable task-specific variants via fine-tuning (Han et al., 2021), reinforcing democratization and innovation (Cupać et al., 2024).
@@ 6. Related Work and Discussion
Relationship analysis between LLM models and datasets. A recent study investigates the practical adaptation of base models to specific tasks. Multitask fine-tuning has demonstrated the potential to enhance performance on target tasks with scarce labels (Xu et al., 2024a). In plant phenotyping, adapting vision-based models by techniques like adapter tuning and decoder tuning has shown results comparable to those of leading task-specific models (Chen et al., 2023). The Quadapter technique for language models tackles quantization difficulties by incorporating learnable parameters that scale activations channel-wise, mitigating overfitting during quantization-aware training (Park et al., 2022).
In future, we would like to explore more through the following perspectives, (i) collecting more LLM supply chain information accurately and scalably; (ii) understanding the LLM supply chain better via both fundamental graph analytics (Feng et al., 2023; Hu et al., 2025; Li et al., 2025) and graph AI techniques (Kipf, 2016; Fu et al., 2022; Wang et al., 2025; Ji, 2025); and (iii) exploring the security and privacy threats on the LLM supply chain (Ji et al., 2021a; He et al., 2024a; Ji et al., 2021b; Cui et al., 2023; Brucker-Hahn et al., 2024).
@@ 7. Conclusion
This project studies the relationships between models and datasets in the LLM ecosystem, which are the central parts of the LLM supply chain. First, we systematically collect the supply chain information of LLMs. With that, we construct a directed heterogeneous graph, having 402,654 nodes and 462,524 edges. Lastly, we perform different types of analysis and make multiple interesting findings.
@@ Acknowledgment
Disclaimer on Generative AI Usage
This work does not use generative AI tools (e.g., ChatGPT, Copilot, Bard, Claude, etc.) for ideation, analysis, code development, or writing. All content, experiments, and results presented in this paper were created and validated solely by the authors. Mentions of generative AI are included only for scholarly context or comparison.